MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2017/10/16
OBSERVER: Paul Rob
CONDITION: Mostly clear
COMMENT: Computers will be shut down in preparation for planned power cut
Compressor gas pressure: 1.75 MPa
Expected Observation Time: START 21:17 END 05:48
Go to the Seeing Log
Go to the Observation Log
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 60583-60593 | 11 | 97 | 5820 |
LMC | 14310-14315 | 6 | 33 | 9900 |
SMC | 1632-1632 | 1 | 2 | 600 |
TRANSIT | - | 0 | 0 | 0 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 18 | 132 | 16320 (04h32m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
4 | 4 | 11 | 10 | 9 | 1 | 2 | 4 | 10 | 9 | 2 | 2 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
4 | 9 | 1 | 2 | 3 | 3 | 1 | 1 | 1 | 1 | 3 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 1 | 1 | 1 | 1 | 6 | 6 | 6 | 2 | 1 | 2 |
Field13 | Field14 | Field15 | Field16 |
2 | 2 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
TRANSIT FIELD STATISTICS
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
21:17 | observing | Partly cloudy |
22:00 | observing | Clear |
23:00 | observing | Clear |
00:00 | observing | Clear |
01:00 | observing | Clear |
02:00 | observing | Mostly clear |
03:00 | observing | Mostly clear |
04:00 | observing | Partly cloudy |
05:00 | waiting | Cloudy |
05:48 | waiting | Cloudy, windy |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day